OpenStudy (anonymous):

cos x + sqrt2= -cos x for x over the interval [0,2pi] I have no idea how to approach this.

OpenStudy (anonymous):

@Mertsj @satellite73 mind seeing if you could help me?

OpenStudy (anonymous):

I know I'm solving for x, but that is all.

OpenStudy (anonymous):

1) add \(\cos(x)\) to both sides 2) subtract \(\sqrt2\) from both sides 3) divide both sides by \(2\) you end up with \[\cos(x)=-\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

that number should look real familiar, find \(x\) so that \(\cos(x)=-\frac{\sqrt2}{2}\)